3x + 5y = -16
3x - 2y = -9

Answer by bb101(1)

3x + 5y = -16 (equation1)
3x-2y = -9 (equation2)
(Equation 1 - Equation2)
7y = -7
(Divide by 7 for both sides)
y= -1
Sub y=-1 into equation 1
3x + 5(-1) = -16
3x -5 = -16
(Subract 5 from both sides)
3x = -11
x = -11/3
therefore x= -11/3 and y=-1